# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

### v3

@@ 5.3 Alternative Mass Measurement
In constraining mass, we also considered a second method for measuring cluster-centric mass elucidated in Remolina González et al. (2021a) using the parametric lens-modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3). For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG. Mass estimates were then obtained from the resulting best-fit lens models according to the methodology described in Remolina González et al. (2021a).
@@
Figure 5: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the parametrically generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work. The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the measured correspondence.
@@
As expected from Remolina González et al. (2021b), these differing methodologies show similar results. The two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ dex as shown in Figure 5. In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster parametric approach throughout the rest of the work.
@@
In theory, lens modelling has the potential to be the most accurate mass-constraining methodology. However, this requires high-spatial-resolution imaging data and ample time. The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands. Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not. These essential positional constraints are often unresolved even with the highest-quality ground-based imaging data. Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions. As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
@@
For more information on lens-modelling software, related techniques, and applications of strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al. (2007), Oguri (2010), Lefor et al. (2013), Meneghetti et al. (2017), Birrer & Amara (2018), Schäfer et al. (2020), Sharon et al. (2022), and Napier et al. (2023a).
@@ 6 Summary and Conclusions
This paper presents measurements of mass and light in 177 strong-lensing systems at cluster and group mass scales. By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically. We use the enclosed $g$, $r$, and $z$-band photometry with Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$. We find that the enclosed total mass and stellar mass for the entire sample of strong-lensing systems considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.443\pm 0.035$, $5.958\pm 0.454$ dex, and $0.173\pm 0.013$ dex respectively. We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$ dex, and $0.161\pm 0.013$ dex respectively. Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon. Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable. The correlations shown in this work should have utility in filtering rankings of possible strong-lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.

### v4

@@ 5.3 Alternative Mass Measurement
In constraining mass, we also considered a second method for measuring cluster-centric mass elucidated in Remolina González et al. (2021a) using the parametric lens-modeling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3). For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (Kassiola & Kovner, 1993) with a free ellipticity locked to the center of the BCG. Mass estimates were then obtained from the resulting best-fit lens models by integrating over the surface-mass-density image, a 2D distribution in the lens plane of the mass fit with LENSTOOL. In this image, each pixel value represents the number of solar masses contained within that pixel. A circular aperture centered on the BCG with area equal to the area enclosed by the best-fit tangential critical curves was used to solve for the LENSTOOL enclosed mass in keeping with Remolina González et al. (2021a).
@@
Figure 7: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the functionally generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work. The black line represents an ideal one-to-one correspondence, and the blue lines represent the uncertainty on the measured correspondence.
@@
As expected from Remolina González et al. (2021b), these differing methodologies show similar results. The two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ dex as shown in Figure 7. In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster functional approach throughout the rest of the work. The agreement between the free-ellipticity-model masses and the circularly symmetric functional masses also demonstrates the feasibility of using a functional mass model for the geometrically simple strong lenses considered in this work.
@@
Lens modeling has the potential to be the most accurate mass-constraining methodology in galaxy clusters. However, this requires high-spatial-resolution imaging data and ample time. The process of iteratively refining just one lens model may take upward of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands. Moreover, the accuracy of measuring mass with lens modeling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not. The essential positional constraints that define these image families are often unresolved—even with the highest-quality ground-based imaging data. Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions. As long as the positional constraints are reasonably placed within the arcs, something that is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
@@
For more information on lens-modeling software, related techniques, and applications strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al. (2007), Oguri (2010), Lefor et al. (2013), Meneghetti et al. (2017), Birrer & Amara (2018), Schäfer et al. (2020), Sharon et al. (2022), and Napier et al. (2023a).
@@ 6 Summary and Conclusions
This paper presents measurements of mass and light in 177 high-confidence cluster-scale strong lenses. By directly obtaining the BCG redshift and an estimate for the Einstein radius for each system, we compute the Einstein-radius-enclosed total mass functionally and independent of any light-traces-mass model. We use the enclosed $g$-, $r$-, and $z$-band LS DR9 photometry with Prospector to estimate the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.148^{+0.051}_{-0.052}X+0.529^{+0.023}_{-0.022}Y+4.121^{+0.272}_{-0.298}$ with an intrinsic scatter on $Z$ of $0.110^{+0.009}_{-0.008}$ dex where $X=z_{{}_{BCG}}$, $Y=\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and $Z=\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$. We find that the enclosed total mass and stellar mass for the entire sample of strong lenses considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.500^{+0.029}_{-0.031}$, $5.254^{+0.395}_{-0.391}$, $0.104^{+0.005}_{-0.003}$ dex, respectively. We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.495^{+0.032}_{-0.033}$, $5.24^{+0.409}_{-0.424}$, $0.103^{+0.005}_{-0.002}$ dex, respectively. We recover scaling relations identical to that of works that measure within cluster-scale apertures—despite the fact that we measure inside apertures with radii some 30–100 times smaller. This is suggestive of invariance in baryon conversion efficiency and feedback strength as a function of cluster-centric radii in galaxy clusters. Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon. Probing these upcoming datasets with both visual and machine-learning-driven searches for new strong lenses at a variety of scales is inevitable. The correlations shown in this work should also have utility in ranking candidate strong lenses as well as offering another measure of cluster cores that can be compared to cosmological simulations.
@@ Acknowledgments
The authors would like to thank the anonymous reviewer for providing detailed feedback, which significantly improved the manuscript. S.D.M. would also like to thank Allison Noble for insightful discussions surrounding the results and discussion Sections of this work.